OpenStudy (johnt):

Determine whether this series is conditionally convergent, absolutely convergent, or divergent

OpenStudy (johnt):

\[\sum_{n = 2}^{\infty} \frac{ 1 }{ \sqrt[3]{n^4 - 1} }\]

OpenStudy (johnt):

Wolfram Alpha tells me it is convergent. I think you should use the comparison test, and try to compare it with a p-series that is convergent. However, I don't know how to identify a p-series that is larger than this series.

OpenStudy (abb0t):

is that a cube root? Try the ratio test. I always use it. lol